OpenStudy (anonymous):

solve for x. (6x-8) - (5x+9)=3

OpenStudy (tux):

x=20

OpenStudy (anonymous):

x=20

OpenStudy (anonymous):

i kind of need to show work on how i got it

OpenStudy (anonymous):

x-17=3

OpenStudy (anonymous):

you combine like terms and then you solve for x.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

whats your number?

OpenStudy (anonymous):

um why??

OpenStudy (anonymous):

to call you and help you

OpenStudy (farmdawgnation):

Ldaniel, asking for other member's phone numbers is completely inappropriate.

OpenStudy (anonymous):

why? i can help better by speaking

OpenStudy (farmdawgnation):

As per the Code of Conduct ( http://openstudy.com/code-of-conduct) we ask that you not solicit other user's personal information for their safety and yours.

OpenStudy (farmdawgnation):

Then perhaps it is best for you to step aside and allow someone else to help?

OpenStudy (farmdawgnation):

@xxguardswaggxx would you still like some help?

OpenStudy (anonymous):

yes, please

OpenStudy (farmdawgnation):

Ok, cool.

OpenStudy (anonymous):

im kind of confused on this question

OpenStudy (farmdawgnation):

Kk. One sec. I'm working it myself to make sure their answer was right before starting with you. heh.

OpenStudy (anonymous):

well thankk you :)

OpenStudy (farmdawgnation):

Alright, so, what we're looking at is this. \[(6x-8)-(5x+9)\] The biggest problem here is we have \(x\) listed twice. What are your thoughts on the best way to fix that?

OpenStudy (anonymous):

that i have to combine both x's so that they're on one side

OpenStudy (farmdawgnation):

Well, they already are on one side, right? But, yes, we do need to combine them.

OpenStudy (farmdawgnation):

So, what can we do to combine them?

OpenStudy (anonymous):

i would think to add 5x

OpenStudy (farmdawgnation):

Not quite. So, here's a hint. We're looking at... \[(6x-8)-(5x+9)=-3\] But we could really consider this to be like... \[(6x-8)-1(5x+9)=-3\] Right? Because that's what the minus sign gives us. So what would we do to get rid of the \(-1\)?

OpenStudy (farmdawgnation):

Sorry. That should have been a positive 3 on the right hand side.

OpenStudy (farmdawgnation):

My B. heh.

OpenStudy (anonymous):

its ok :)

OpenStudy (farmdawgnation):

So, \((6x−8)−1(5x+9)=3\) But yeah, how would we get rid of this -1?

OpenStudy (farmdawgnation):

Errr. I'm about to have to leave. :\

OpenStudy (anonymous):

sorry, im getting it

OpenStudy (anonymous):

you would add the negative 1

OpenStudy (farmdawgnation):

Err, not quite. The negative one is attached to the \((5x+9)\). That piece could also be written at \(-1\times(5x+0)\), right?

OpenStudy (farmdawgnation):

Er, heh. Typing fail. \(−1\times(5x+9)\)

OpenStudy (anonymous):

ahh yeah so it could be written as: -5x+(-9)

OpenStudy (farmdawgnation):

Indeed. Or, better yet, \(-5x-9\).

OpenStudy (anonymous):

yeah

OpenStudy (farmdawgnation):

So, what does that give us in the overall equation?

OpenStudy (anonymous):

6x-8 -5x-9=3

OpenStudy (farmdawgnation):

You got it, boss. Now combine terms.

OpenStudy (farmdawgnation):

Should be smooth sailing after that.

OpenStudy (anonymous):

haha yeah c: thanks

OpenStudy (farmdawgnation):

Np. Happy to help. Best of luck! I'm outta here for now.

OpenStudy (anonymous):

alright see ya!